# Collecting Cost-Effective, High-Quality Truthfulness Assessments with LLM Summarized Evidence

### v1

@@ 3. RQ3:
In evaluating the practical aspects of crowdsourced truthfulness assessments, time efficiency emerges as a critical metric, as the more time a worker spends on the task the higher the required cost for its judgment and, as consequence, for the overall task. Figure 5 shows the time required for workers to provide assessments for both task modalities. As shown by inspecting the two series in the plot, the average time taken to complete assessments consistently decreases as the statement index increases (i.e., as a worker performs the task). This trend suggests that participants become more proficient at navigating and processing the tasks over time. Such a learning effect, also observed in previous studies [25, 48], suggests the adaptability of participants and their increasing efficiency in handling the assessments, whether they are dealing with full-length or summarized content.
When comparing the two series we see that the Summary modality consistently shows a reduced average time on each statement index compared to the Standard modality, with the only exception being the second document, and the difference is statistically significant according to the Mann-Whitney U Test [31] at the $p<0.05$ level. This supports the observation that time efficiency gained from presenting workers with summarized evidences are real. In practice, the reduced time spent in the Summary modality can lead to enhanced productivity and a lighter cognitive load for participants, which can perform more judgment in the same amount of time and are required to read much less text. Such improvements in efficiency could also translate into cost savings for researchers and practitioners implementing these misinformation assessments at scale, potentially enhancing the economic feasibility of deploying crowdsourced truthfulness evaluations. This can be seen by inspecting the cumulative series in Figure 5, which show the potential for cost savings in large-scale truthfulness assessments. For instance, completing 600 assessments, equivalent to 120 documents, requires approximately 23.67 hours of total work and costing £171.60 at the U.S. minimum wage of \$7.25 per hour (approximately £6.08). In contrast, the Summary modality, reduces the required total work to approximately 20.50 hours, costing £148.63 under the same wage assumption. While these figures are based on the minimum wage, actual costs would scale proportionally if higher wages are offered. When extended to larger datasets requiring (let us say) 6,000 assessments, the Standard modality incurs costs of £1,716, compared to £1,486 for the Summary modality, resulting in higher savings. These findings emphasize the economic advantages of the Summary modality, which maintains comparable performance metrics while significantly reducing worker time and overall costs, making it particularly valuable for large-scale annotation efforts.
@@
Previously, we used the Standard modality as the baseline, stacking judgments in the Summary modality to match the total time spent on assessments in the Standard batch. In this analysis, we reverse the approach by sampling judgments from the Standard modality to match the total time spent in the Summary batch, using the Summary modality as the baseline. We do that by using a simulation-based analysis. In each of the 1,000 simulations, we randomly sample assessments from the Standard modality to match the total time spent in the Summary modality. The results are shown in Figure 6. The average effectiveness scores achieved by the Standard modality across the simulations indicates that the Standard modality maintains a comparable level of accuracy even when constrained by the same amount of time. Also in this case no statistical significance emerges between the two modalities. These findings suggest that both modalities perform similarly in terms of absolute deviations from the ground truth, with no significant disparities in judgment quality.
@@
In the context of crowdsourced truthfulness assessments, understanding the perceived usefulness of the evidence is important for understanding the quality and trustworthiness of the workers’ provided assessments. This provides insight into how the evidence impacts the decision-making process and the reliance of participants on the provided information. Figure 7 shows the perceived usefulness and necessity of the evidence for the two modalities. We start with inspecting the need for evidence as expressed by participants across different truthfulness levels. Data shows a significant reliance on evidence, with the majority of responses across both standard and summary modalities suggesting that assessments are “fully based on evidence”. This strong dependency highlights the critical role that evidence plays in ensuring accurate and reliable assessments. Notably, in the Summary modality, there is a slightly higher tendency for assessments to be fully based on evidence compared to the Standard modality, suggesting that even summarized content sufficiently supports robust judgment formation. Less frequently, participants report that their assessments were “more than my initial opinion” or “considered initial opinion”, indicating that while evidence guides their evaluations, preconceived notions or initial impressions also play a role. Interestingly, the number of participants who report their assessments as “considered slightly” or “did not consider” the evidence is minimal, suggesting the minimal influence of unsupported personal bias in the evaluation process.
Turning to the usefulness of evidence, participants across both modalities rate the majority of the evidence as being “extremely useful” or “very useful”. This high level of perceived usefulness is consistent across all truthfulness categories, remarking the importance of quality evidence in the assessment task. In the Summary modality, there is a notable pattern where evidence is frequently considered at least “moderately useful”, which aligns with the reduced amount of information presented to the workers, yet keeping its impact on their assessments. The less frequent ratings of “slightly useful” or “Not useful at all” are significantly lower, which indicates that the evidence provided is overall useful to workers. This is true also for the Summary modality, where even summarized information appears to be adequately informative to workers.
@@
We now detail the result of the inspection of the number of evidences participants interacted with in both the standard and summary modalities during the truthfulness assessment tasks. We focus particularly on the number of evidences clicked by participants for each statement. A specific count of interest was 8, which indicates that participants clicked on exactly one link per statement (including gold questions), suggesting a uniform and minimal engagement across all statements provided in the task. For the Summary modality, we found that out of the total interactions, 511 instances (85.17%) involved clicking exactly eight links to evidence. Similarly, in the Standard modality, there were 520 instances (86.67%) where participants engaged with exactly eight links. This close similarity in percentages between the two modalities indicates a comparable level of engagement when workers are provided with either full or summarized content, in the case of workers that do the minimum amount of work. A similar pattern emerges when we exclude instances where participants clicked on exactly eight links. In such cases, the mean number of links clicked in the Summary modality was $18.70$, with a standard deviation of $8.09$, whereas in the Standard modality, the mean was slightly lower at $17.50$, with a narrower standard deviation of $3.39$. This is based on few (approx 15%) of instances, and no statistical significance is measured.
@@
Our findings show that in the Summary modality, 433 instances (72.17%) involved participants clicking on each dimension only once, thus doing the minimum required number of clicks. Similarly, in the Standard modality, there were 450 instances (75%) where participants also clicked each dimension just once. When we exclude cases where the number of clicks on dimensions is exactly three, we get that for these cases, the Summary modality shows a mean of $4.45$ clicks with a standard deviation of $0.67$, slightly less than the Standard modality, which shows a mean of $4.60$ clicks and a standard deviation of $0.99$. No statistical significance is detected. These statistics suggest that participants in both modalities generally stick to their initial assessments. The similar levels of engagement suggest that summarized content appears to be sufficient for making assessments. The slightly lower variance in the number of clicks in the Summary modality further reinforces this point, suggesting that participants are slightly more consistently confident in their assessments.
@@
To further explore the interplay between key dimensions in the assessment process, we conduct a correlation analysis examining the relationships among need, usefulness, and time elapsed for both modalities (Table 5). The results reveal a consistent moderate positive correlation ($\rho\approx 0.48$) between need and usefulness, indicating that participants who perceive the evidence as more necessary also tend to consider it more useful, irrespective of whether they are presented with full-length or summarized content. This relationship underscores the complementary nature of these two dimensions, as both contribute to the assessment process. However, the moderate strength of the correlation suggests that while need and usefulness are related, they are not interchangeable. Instead, these dimensions capture distinct aspects of participants’ interactions with evidence. On the other hand, the correlations between need and time elapsed, as well as usefulness and time elapsed, are negligible or weak ($\rho\approx-0.02$ to $-0.09$), showing limited influence of these perceptions on task duration.
In order to more deeply analyze interactions between the three assessment dimensions measured for the two modalities (i.e., time elapsed, need, and usefulness), we performed an OLS regression analysis, augmented by $\omega^{2}$ effect sizes [37], which is well-established in the literature for quantifying effect sizes [17, 54, 18, 42, 16]. The results confirm that the Summary modality significantly reduces task time when compared to the Standard modality ($p<0.05$, $\omega^{2}=0.0036$). Conversely, perceptions of need, usefulness, and their interaction exhibited no statistically significant effects on task duration, with negligible $\omega^{2}$ values ($-0.0008$ for need, $-0.0001$ for usefulness, and $-0.0005$ for their interaction).
@@
Regarding RQ2 (Efficiency), we find that the Summary modality substantially reduces the time required for participants to complete their assessments compared to the Standard one. Workers using summarized evidence complete nearly double the number of judgments within the same time frame. This efficiency gain highlights the potential of summarization to enhance throughput in large-scale truthfulness assessment tasks, particularly in time-sensitive scenarios. Addressing cost-effectiveness, we find that summarization significantly reduces the costs associated with large-scale truthfulness assessments. By enabling workers to complete more judgments within the same time frame, the Summary modality reduces overall task costs without sacrificing quality. This economic advantage is particularly relevant for scaling fact-checking efforts where budget constraints are a key consideration.
In terms of RQ3 (Behavioral Insights), we observe that the Summary modality does not diminish participants’ reliance on or perceived usefulness of evidence. Furthermore, the Summary modality produces higher internal agreement among workers, suggesting that the condensed presentation of information aligns participant judgments more effectively. This indicates that summarization not only maintains but also enhances certain aspects of the decision-making process, supporting consistent and evidence-based assessments.
Despite these contributions, certain limitations must be acknowledged. First, the reliance on automatically generated summaries introduces the possibility of omitting critical nuances present in full-length evidence, which may impact worker judgment in complex cases. Additionally, our study focuses on a specific dataset and task context, which may limit the generalizability of our findings to other domains or settings. Finally, while the Summary modality demonstrates promising efficiency, further research is needed to explore its long-term effects on worker fatigue, engagement, and the potential biases introduced by condensed evidence.
While these findings highlight the promise of summarization, they also leave analysis open for future exploration. One direction is to refine summarization techniques to better tailor evidence to specific truthfulness categories or individual worker preferences, which could further improve accuracy and efficiency. Another area of investigation is the impact of evidence presentation on worker satisfaction and engagement, which could offer insights into designing more intuitive and rewarding tasks. Additionally, applying these methods to other domains, such as health misinformation or scientific claims, would test their applicability and effectiveness in different contexts. Another promising direction for future work is the exploration of multi-document summarization techniques. By consolidating multiple pieces of evidence into a single, coherent summary, this approach could potentially improve worker accuracy and reduce cognitive load, especially when participants prefer to engage with a single document.
@@ Acknowledgments
[Uncaptioned image][Uncaptioned $\end{array}$ This research is partially funded by the European Union (ERC, ExplainYourself, 101077481), as well as a Danish Data Science Academy postdoctoral fellowship (grant number: 2023-1425) and is supported by the Pioneer Centre for AI, DNRF grant number P1.
This research is partially supported by the PRIN 2022 Project – “MoT—The Measure of Truth: An Evaluation-Centered Machine-Human Hybrid Framework for Assessing Information Truthfulness” (Code No. 20227F2ZN3, CUP No. G53D23002800006), funded by the European Union – Next Generation EU – PNRR M4 C2 I1.1.

### v2

@@ 3. RQ3:
In evaluating the practical aspects of crowdsourced truthfulness assessments, time efficiency emerges as a critical metric, as the more time a worker spends on the task, the higher the cost required for their judgment and, consequently, for the overall task. Figure 5 shows the time required for workers to provide assessments for both task modalities. As shown by inspecting the two series in the plot, the average time taken to complete assessments consistently decreases as the statement index increases (i.e., as a worker performs the task). This trend suggests that participants become more proficient at navigating and processing the tasks over time. Such a learning effect, also observed in previous studies [28, 51], suggests the adaptability of participants and their increasing efficiency in handling the assessments, whether they are dealing with full-length or summarized content.
When comparing the two series we see that the Summary modality consistently shows a reduced average time on each statement index compared to the Standard modality, with the only exception being the second document, and the difference is statistically significant according to the Mann-Whitney U Test [34] at the $p<0.05$ level. This supports the observation that time efficiency gained from presenting workers with summarized evidence are real. In practice, the reduced time spent in the Summary modality can lead to enhanced productivity and a lighter cognitive load for participants, which can perform more judgment in the same amount of time and are required to read much less text. Such improvements in efficiency could also translate into cost savings for researchers and practitioners implementing these misinformation assessments at scale, potentially enhancing the economic feasibility of deploying crowdsourced truthfulness evaluations. This can be seen by inspecting the cumulative series in Figure 5, which show the potential for cost savings in large-scale truthfulness assessments.
For instance, completing 600 assessments, equivalent to 120 documents, requires approximately 23.67 hours of total work and costing £171.60 at the U.S. minimum wage of \$7.25 per hour (approximately £6.08). In contrast, the Summary modality, reduces the required total work to approximately 20.50 hours, costing £148.63 under the same wage assumption. While these figures are based on the minimum wage, actual costs would scale proportionally if higher wages are offered. When extended to larger datasets requiring (let us say) 6,000 assessments, the Standard modality incurs costs of £1,716, compared to £1,486 for the Summary modality, resulting in higher savings. These findings emphasize the economic advantages of the Summary modality, which maintains comparable performance metrics while significantly reducing worker time and overall costs, making it particularly valuable for large-scale annotation efforts.
@@
Previously, we used the Standard modality as the baseline, stacking judgments in the Summary modality to match the total time spent on assessments in the Standard batch. In this analysis, we reverse the approach by sampling judgments from the Standard modality to match the total time spent in the Summary batch, using the Summary modality as the baseline. We do that by using a simulation-based analysis. In each of the 1,000 simulations, we randomly sample assessments from the Standard modality to match the total time spent in the Summary modality. The results are shown in Figure 6. The average effectiveness scores achieved by the Standard modality across the simulations indicates that the Standard modality maintains a comparable level of accuracy even when constrained by the same amount of time. As in the previous case, no statistical significance emerges between the two modalities. These findings suggest that both modalities perform similarly in terms of absolute deviations from the ground truth, with no significant disparities in judgment quality.
@@
In the context of crowdsourced truthfulness assessments, understanding the perceived usefulness of the evidence is important for understanding the quality and trustworthiness of the workers’ provided assessments. This provides insight into how the evidence impacts the decision-making process and the reliance of participants on the provided information. Figure 7 shows the perceived usefulness and necessity of the evidence for the two modalities.
We start with inspecting the need for evidence as expressed by participants across different truthfulness levels. Data shows a significant reliance on evidence, with the majority of responses across both Standard and Summary modalities suggesting that assessments are “fully based on evidence”. This strong dependency highlights the critical role that evidence plays in ensuring accurate and reliable assessments. Notably, in the summary modality, there is a slightly higher tendency for assessments to be fully based on evidence compared to the standard modality, suggesting that even summarized content sufficiently supports robust judgment formation. Less frequently, participants report that their assessments were “more than my initial opinion” or “considered initial opinion”, indicating that while evidence guides their evaluations, preconceived notions or initial impressions also play a role. Interestingly, the number of participants who report their assessments as “considered slightly” or “did not consider” the evidence is minimal, suggesting the minimal influence of unsupported personal bias in the evaluation process.
Turning to the usefulness of evidence, participants across both modalities rate the majority of the evidence as being “extremely useful” or “very useful”. This high level of perceived usefulness is consistent across all truthfulness categories, remarking the importance of quality evidence in the assessment task. In the Standard modality, there is a notable pattern where evidence is frequently considered at least “moderately useful”, which aligns with the reduced amount of information presented to the workers, yet keeping its impact on their assessments. The less frequent ratings of “slightly useful” or “not useful at all” are significantly lower, which indicates that the evidence provided is overall useful to workers. This holds as well for the Summary modality: also summarized information appears to be adequately informative to workers.
@@
We now detail the result of the inspection of the number of evidences participants interacted with in both the standard and summary modalities during the truthfulness assessment tasks. We focus particularly on the number of evidences clicked by participants for each statement. A specific count of interest was 8, which indicates that participants clicked on exactly one link per statement (including gold questions), suggesting a uniform and minimal engagement across all statements provided in the task. For the Summary modality, we found that out of the total interactions, 511 instances (85.17%) involved clicking exactly eight links to evidence. Similarly, in the Standard modality, there were 520 instances (86.67%) where participants engaged with exactly eight links. This close similarity in percentages between the two modalities indicates a comparable level of engagement when workers are provided with either full or summarized content, in the case of workers who do the minimum amount of work. A similar pattern emerges when we exclude instances where participants clicked on exactly eight links. In such cases, the mean number of links clicked in the Summary modality was $18.70$, with a standard deviation of $8.09$, whereas in the Standard modality, the mean was slightly lower at $17.50$, with a narrower standard deviation of $3.39$. This is based on few (approx 15%) of instances, and no statistical significance is measured.
@@
Our findings show that in the Summary modality, 433 (72.17%) of participants clicked each dimension only once, similar to 450 (75%) in the Standard modality. Excluding cases with exactly three clicks, the Summary modality shows a mean of $4.45$ clicks ($\sigma^{2}=0.67$), slightly lower than the Standard modality’s $4.60$ clicks ($\sigma^{2}=0.99$). No statistical significance is detected. These statistics suggest that participants in both modalities generally stick to their initial assessments, and the similar levels of engagement suggest that summarized content appears to be sufficient for making assessments. The slightly lower variance in the number of clicks in the Summary modality further reinforces this point, suggesting that participants are slightly more consistently confident in their assessments.
@@
To further explore the interplay between these dimensions, we conduct a correlation analysis examining the relationships among need, usefulness, and time elapsed for both modalities (Table 5). The results reveal a consistent moderate positive correlation ($\rho\approx 0.48$) between need and usefulness, indicating that participants who perceive the evidence as more necessary also tend to consider it more useful, irrespective of whether they are presented with full-length or summarized content. This relationship underscores the complementary nature of these two dimensions, as both contribute to the assessment process. However, the moderate strength of the correlation suggests that while need and usefulness are related, they are not interchangeable. Instead, these dimensions capture distinct aspects of participants’ interactions with evidence. On the other hand, the correlations between need and time elapsed, and usefulness and time elapsed, are negligible ($\rho\approx-0.02$ to $-0.09$), showing limited influence of these perceptions on task duration.
To more deeply analyze the interactions among the three assessment dimensions measured for the two modalities (i.e., time elapsed, need, and usefulness), we performed an OLS regression analysis, augmented with $\omega^{2}$ effect sizes [40], a well-established method in the literature for quantifying effect sizes [20, 61, 21, 45, 19]. The results confirm that the Summary modality significantly reduces task time when compared to the Standard modality ($p<0.05$, $\omega^{2}=0.0036$). Conversely, perceptions of need, usefulness, and their interaction exhibited no statistically significant effects on task duration, with negligible $\omega^{2}$ values ($-0.0008$ for need, $-0.0001$ for usefulness, and $-0.0005$ for their interaction).
@@
Regarding RQ2 (Efficiency), we find that the Summary modality substantially reduces the time required for participants to complete their assessments compared to the Standard one. Workers using summarized evidence complete nearly double the number of judgments within the same time frame. Addressing cost-effectiveness, we find that summarization significantly reduces the costs associated with large-scale truthfulness assessments. By enabling workers to complete more judgments within the same time frame, the Summary modality reduces overall task costs without reducing quality.
In terms of RQ3 (Behavioral Insights), we observe that the Summary modality does not diminish participants’ reliance on or perceived usefulness of evidence. Furthermore, the Summary modality produces higher internal agreement among workers, suggesting that the condensed presentation of information aligns participant judgments more effectively. This indicates that summarization not only maintains but also enhances certain aspects of the decision-making process, supporting consistent evidence-based assessments.
Despite these contributions, certain limitations must be acknowledged. First, the reliance on automatically generated summaries introduces the possibility of omitting critical nuances present in full-length evidence, which may impact worker judgment in complex cases. Additionally, our study focuses on a specific dataset and task context, which may limit the generalizability of our findings. Finally, while the Summary modality demonstrates promising efficiency, further research is needed to explore its long-term effects on worker fatigue, engagement, and the potential biases.
These findings also suggest directions for future work. We plan to refine summarization techniques to better tailor evidence to specific truthfulness categories or individual worker preferences; also, we plan to test the impact of evidence presentation on worker satisfaction and engagement. Additionally, we plan to generalize these methods to other domains, such as health misinformation or scientific claims. Finally, we plan to explore multi-document summarization techniques that would allow to present multiple pieces of evidence into a single, coherent summary.
@@ Acknowledgments
This research is partially funded by the European Union (ERC, ExplainYourself, 101077481), as well as a Danish Data Science Academy postdoctoral fellowship (grant number: 2023-1425) and is supported by the Pioneer Centre for AI, DNRF grant number P1. It is also partially supported by the PRIN 2022 Project, “MoT—The Measure of Truth: An Evaluation-Centered Machine-Human Hybrid Framework for Assessing Information Truthfulness”, Code No. 20227F2ZN3, CUP No. G53D23002800006, funded by the European Union – Next Generation EU – PNRR M4 C2 I1.1.